graph the equation by plotting points
y=x+2

Answer by ewatrrr(24785)

 
Hi,
graph the equation by plotting points
y=x+2
y = 0 + 2
0 = x + 2
Pt(0,2) and Pt(-2,0) on this line
Plotting points and connecting with a Line